OpenStudy (anonymous):

use l'hopital's rule to evaluate the limit x^5-1/x-1 as x approaches 1

OpenStudy (anonymous):

l'Hopital's rule just says take the derivative of top and bottom of the fraction and the limit will be the same, assuming that it is an indeterminate form... which we can assume is the case if they tell you to use the rule.

OpenStudy (anonymous):

So plug \(1\) into \[\frac{\frac{d}{dx}x^5-1}{\frac{d}{dx}x-1}\]